Two volatile liquids A and B have vapour pressures of 222 Torr and 444 Torr respectively in an ideal binary solution at 80°C. What is the vapour pressures of pure liquid A if vapour pressure of pure liquid B is 555 Torr, at 80°C?

Grade:12

2 Answers

Priyanshu Gujjar
2 years ago

Total vapour pressure P=PA+PB

PA=100×2+32=100×52=40 torr
PB=80×2+33=80×53=48 torr
PTotal=PA+PB=40+48=88 t
Ansh Shah
2 years ago
PA=222
PB=444
PoB=555
PoA=?
Pb=Pob*xb
hence, xb=0.8
we know that for a binary solution, xa+xb=1
so xa=1-0.8=1-Xb=0.2
so,
poa=pa/xa=222/0.2=1110 Torr.